A open box is to be made from a 25inch by 40inch rectangular piece of cardboard by cutting squares of equal area from each corner and turning up the sides. If the specifications require the base the base of the box to have an area of 700sq.in., what size squares should be cut from each corner?
(Explain what the answer means in mathematical perspective.)

Answer by checkley77(12844)

(25-2x)(40-2x)=700
1,000-130x+4x^2=700
4x^2-130x+1,000-700=0
4x^2-130x+300=0
(2x-60)(2x-5)=0
2x-5=0
2x=5
x=5/2
x=2.5 in. is the sqare that is cut.
Proof:
(25-2*2.5)(40-2*2.5)=700
(25-5)(40-5)=700
20*35=700
700=700